The sum of three consecutive natural numbers is 963, find the numbers.
(A) First, write out an equation that models the relationship between these three numbers:
x + + = 963
2nd number 3rd number

(B) Now, simplify the left side of your equation:
= 963
(C) Solve your equation and identify the three numbers:
The three numbers in increasing order are , and .


Answer by Timnewman(323)

Hello dear,
Do not post question that you know it's solution,rather post the one that looks a bit technical to you.
incase you do not know this one,follow the steps below.
Let the three consecutive nembers be x,x+1,x+2.
From your question,
x+x+1+x+2=963
3x+3=963
3x=963-3
3x=960
x=320.
The number's are 320,321,322
Was this helpfull?